40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 10 19:56:23 2009, rana, koji, Omnistructure, Environment, Changed office temp therm.JPG
    Reply  Sun Jul 12 20:05:06 2009, rana, Omnistructure, Environment, Changed office temp Picture_3.png
       Reply  Mon Jul 13 15:48:05 2009, rana, Omnistructure, Environment, Removal of the cold air deflection device for the MOPA chiller 
          Reply  Mon Jul 13 23:03:14 2009, rob, alberto, Omnistructure, Environment, Removal of the cold air deflection device for the MOPA chiller 
             Reply  Tue Jul 14 00:32:46 2009, rob, alberto, Omnistructure, Environment, Removal of the cold air deflection device for the MOPA chiller 
                Reply  Tue Jul 14 17:48:20 2009, Jenne, Omnistructure, Environment, Removal of the cold air deflection device for the MOPA chiller ACeffectonLASER2.png
Message ID: 1745     Entry time: Tue Jul 14 17:48:20 2009     In reply to: 1741
Author: Jenne 
Type: Omnistructure 
Category: Environment 
Subject: Removal of the cold air deflection device for the MOPA chiller 

Quote:

Quote:

Quote:
Around 2 PM today, I removed the blue flap which has been deflecting the cold air from the AC down into the laser chiller.
Let's watch the laser trends for a few days to see if there's any effect.


Alberto has moved us to stage 2 of this experiment: turning off the AC.

The situation at the control room computers with the AC on minus the blue flap is untenable--it's too cold and the air flow has an unpleasant eye-drying effect.


I turned the AC back on because the temperature of the room was going up so also that of the laser chiller.


I reinstalled the blue-flap technology on the AC, because the MOPA power was dropping like a rock. A light-ish rock since it wasn't going down too fast, but the alarms started going a little while ago because PMC trans was too low, because the power was getting a little low. The laser water chiller is reading 21.97C, which is higher than it normally does/did before the AC shenanigans (It usually reads 20.00C).

Attached is a look-back of 18 hours, during which you can see in the AMPMON the time that Rana removed the blue flap around 2pm yesterday and the AMPMON changes a little bit, but not drastically, the time around 11pm when the AC was turned off, and AMPMON goes down pretty fast, and about 12:30am, when Alberto turned the AC back on, and AMPMON starts to recover. I think that the AMPMON starts to go down again in the morning because it's been crazy hot here in Pasadena, so the room might be getting warmer, especially with the laser chiller-chiller not actively chilling the laser chiller (by not being pointed at the water chiller), so the water isn't getting as cold, and the HTEMP started to go up.

In the last few minutes of having put the blue flap back on the AC, the laser chiller is already reading a lower temperature, and the AMPMON is starting to recover.
Attachment 1: ACeffectonLASER2.png  18 kB  | Hide | Hide all
ACeffectonLASER2.png
ELOG V3.1.3-